Consider the following sample of exam scores: 

x = {28,57,58,64,79,80,83,85,87,89,89,90, 94,94}

Obtain the sample standard deviation.



1
Expert's answer
2021-07-27T16:38:35-0400

"\bar x=\frac{1}{14}\Sigma_{i=1}^{14}x_i=\frac{1077}{14}=76.9286."

"s=\sqrt{\frac{1}{13}\Sigma_{i=1}^{14}(x_i-\bar x)^2}=\sqrt{\frac{4578.9286}{13}}=\sqrt{352.2253}=18.7677."